OpenStudy (anonymous):

^3√3x-1=2 show all work please

OpenStudy (johnweldon1993):

okay so you have \[\sqrt[3]{3x} - 1 = 2\] or is it \[\sqrt[3]{3x-1} = 2\]?

OpenStudy (anonymous):

the second one .